# SeeThruFinger: See and Grasp Anything with a Soft Touch

### v1

@@ Abstract
We present SeeThruFinger, a soft robotic finger with an in-finger vision for multi-modal perception, including visual perception and tactile sensing, for geometrically adaptive and real-time reactive grasping. Multi-modal perception of intrinsic and extrinsic interactions is critical in building intelligent robots that learn. Instead of adding various sensors for different modalities, a preferred solution is to integrate them into one elegant and coherent design, which is a challenging task. This study leverages the Soft Polyhedral Network design as a robotic finger, capable of omni-directional adaptation with an unobstructed view of the finger’s spatial deformation from the inside. By embedding a miniature camera underneath, we achieve the visual perception of the external environment by inpainting the finger mask using E2FGV, which can be used for object detection in the downstream tasks for grasping. After contacting the objects, we use real-time object segmentation algorithms, such as XMem, to track the soft finger’s spatial deformations. We also learned a Supervised Variational Autoencoder to enable tactile sensing of 6D forces and torques for reactive grasp. As a result, we achieved multi-modal perception, including visual perception and tactile sensing, and soft, adaptive object grasping within a single vision-based soft finger design compatible with multi-fingered robotic grippers.
@@ 3.2 The SeeThruFinger Architecture with Robot Learning
This study proposes the SeeThruFinger architecture for learning visual perception and tactile sensing simultaneously via the in-finger vision with a soft touch, as shown in Fig. 1. Besides omni-directional adaptation for grasping, the SeeThruFinger also features an in-finger vision with direct line-of-sight to the external environment but suffers from occlusions by the finger network. By segmenting the finger mask either through simple thresholding or with clickable interaction with SAM [3], we implement the inpainting algorithm, i.e., E2FGV [30], to generate the occluded scene and objects for visual perception. Then, we performed object detection using the inpainted image, i.e., Real-Time Detection Transformer (RT-DETR) [31], generating object class and its planar coordinates $(x,y)$ on the tabletop for manipulation planning. During its first attempt, the gripper will attempt to pick up the object at a fixed pose and depth $z_{0}$ relative to the robot. While closing the fingers, the same SeeThruFinger detects if it has contact with the object by inferring tactile feedback in 6D forces and torques based on the finger’s whole-body deformation. Usually, contact normal to the finger surface will cause the finger to deform adaptively for enhanced grasping results [32]. Due to the network design, the finger can deform twistedly to generate torque about the $z$ axis. When $\tau_{z}$ is detected to be larger than 0.05 Nm, the object orientation is severely different from the gripper’s initial pose, suggesting a less ideal grasping that can be optimized through regrasping by turning the wrist joint reversely to reduce $\tau_{z}$ for an enhanced adaptation and grasping robustness. We implemented a Supervised Variational Autoencoder (SVAE) to enable tactile sensing using SeeThruFinger’s design by tracking the finger’s whole-body deformation after contact.
@@ 4.1 In-Finger Scene Inpainting and Object Detection with Static Occlusion
For inpainting, we first obtain the mask template $\mathbf{M_{0}}$ in Fig. 3(d) using threshold segmentation against a clean background. Results show that the soft finger occludes about 53.6% of the image pixels. Alternatively, we can use SAM to obtain the mask by clicking for comparison in an actual scene. Fig. 3(e) shows the inpainting operation implemented with E2FGV [30] using raw image stream from in-finger vision and the finger mask from threshold segmentation, generating the scene with objects on the tabletop. The inpainted images managed to reconstruct sufficient details of the occluded objects and the scene, as shown in the dashed white boxes in Fig. 3(c). We tested the feasibility of the inpainted scene by performing object detection using RT-DETR [31] based on the inpainted scene, producing object classification (%) and bounding box center ($x,y$) for manipulation planning. Fig. 3(f) is an example of the object detection performance, which is acceptable for simple grasping. Although it would be interesting to test with other more advanced pose estimation algorithms, we intentionally used a simpler algorithm that generates only the center of the bounding box to test reactive grasping with soft adaptation [32] and tactile sensing [33] using the SeeThruFinger.

### v2

# SeeThruFinger: See and Grasp Anything with a Multi-Modal Soft Touch

@@ Abstract
We present SeeThruFinger, a Vision-Based Tactile Sensing (VBTS) architecture using a markerless See-Thru-Network. It achieves simultaneous visual perception and tactile sensing while providing omni-directional, adaptive grasping for manipulation. Multi-modal perception of intrinsic and extrinsic interactions is critical in building intelligent robots that learn. Instead of adding various sensors for different modalities, a preferred solution is to integrate them into one elegant and coherent design, which is a challenging task. This study leverages the in-finger vision to inpaint occluded regions of the external environment, achieving coherent scene reconstruction for visual perception. By tracking real-time segmentation of the Soft Polyhedral Network’s large-scale deformation, we achieved real-time markerless tactile sensing of 6D forces and torques. We further demonstrate the application of the SeeThruFinger for reactive grasping without using external cameras or dedicated force and torque sensors. As a result, our proposed SeeThruFinger architecture enables multi-modal perception via a single in-finger vision camera in a markerless way, including scene inpainting, object detection, segmentation tracking, and tactile sensing.
@@ 3.2 The SeeThruFinger Architecture with Robot Learning
This study proposes the SeeThruFinger architecture for learning visual perception and tactile sensing simultaneously via the in-finger vision with a soft touch, as shown in Fig. 1. Besides omni-directional adaptation for grasping, the SeeThruFinger also features an in-finger vision with direct line-of-sight to the external environment but suffers from occlusions by the finger network. By segmenting the finger mask either through simple thresholding or with clickable interaction with SAM [3], we implement the inpainting algorithm, i.e., E2FGVI [30], to generate the occluded scene and objects for visual perception. Then, we performed object detection using the inpainted image, i.e., Real-Time Detection Transformer (RT-DETR) [31], generating object class and its planar coordinates $(x,y)$ on the tabletop for manipulation planning. During its first attempt, the gripper will attempt to pick up the object at a fixed pose and depth $z_{0}$ relative to the robot. While closing the fingers, the same SeeThruFinger detects if it has contact with the object by inferring tactile feedback in 6D forces and torques based on the finger’s whole-body deformation. Usually, contact normal to the finger surface will cause the finger to deform adaptively for enhanced grasping results [32]. Due to the network design, the finger can deform twistedly to generate torque about the $z$ axis. When $\tau_{z}$ is detected to be larger than 0.05 Nm, the object orientation is severely different from the gripper’s initial pose, suggesting a less ideal grasping that can be optimized through regrasping by turning the wrist joint reversely to reduce $\tau_{z}$ for an enhanced adaptation and grasping robustness. We implemented a Supervised Variational Autoencoder (SVAE) to enable tactile sensing using SeeThruFinger’s design by tracking the finger’s whole-body deformation after contact.
@@ 4.1 In-Finger Scene Inpainting and Object Detection with Static Occlusion
For inpainting, we first obtain the mask template $\mathbf{M_{0}}$ in Fig. 3(d) using threshold segmentation against a clean background. Results show that the soft finger occludes about 53.6% of the image pixels. Alternatively, we can use SAM to obtain the mask by clicking for comparison in an actual scene. Fig. 3(e) shows the inpainting operation implemented with E2FGVI [30] using raw image stream from in-finger vision and the finger mask from threshold segmentation, generating the scene with objects on the tabletop. The inpainted images managed to reconstruct sufficient details of the occluded objects and the scene, as shown in the dashed white boxes in Fig. 3(c). We tested the feasibility of the inpainted scene by performing object detection using RT-DETR [31] based on the inpainted scene, producing object classification (%) and bounding box center ($x,y$) for manipulation planning. Fig. 3(f) is an example of the object detection performance, which is acceptable for simple grasping. Although it would be interesting to test with other more advanced pose estimation algorithms, we intentionally used a simpler algorithm that generates only the center of the bounding box to test reactive grasping with soft adaptation [32] and tactile sensing [33] using the SeeThruFinger.